# Exploring Communication Dynamics: Eye-tracking Analysis in Pair Programming of Computer Science Education

### v1

@@ 4.1. Eye-Tracking Analyses
We performed the Shapiro-Wilk test to verify the normality assumption. If it did not meet the normality assumption, the Mann-Whitney U test was used for independent samples  (Kim, 2015) and the Kruskal-Wallis (Kruskal and Wallis, 1952) was used to assess the differences over two groups as a nonparametric statistical test. Thus, the Mann-Whitney U test revealed significant differences in the total duration of whole fixations between positions (U = 0.533) and roles (U = 0.530), respectively. Similarly, the number of whole fixations exhibited considerable differences between positions (U = 0.526) and between roles (U = 0.531). Moreover, notable differences were observed in the number of saccades in various positions (U = 0.536), roles (U = 0.534), and groups (KW = 6.346). All p-values were significant (p ¡ .001) in this result, excluding the number of saccades in groups (p ¡ .05).
@@ 4.2. Questionnaires & FGI
Expert→Expert 13 3.73 0.45
Expert→Student 6 3.39 0.61
Student→Expert 9 3.90 0.22
Student→Student 8 3.62 0.49

### v3

@@ 4.1. Eye-Tracking Analyses
We performed the Shapiro-Wilk test to verify the normality assumption. As the data did not meet the normality assumption, the Mann-Whitney U test was used for independent samples  (Kim, 2015), and the Kruskal-Wallis (Kruskal and Wallis, 1952) was used to assess the differences among three groups as a nonparametric statistical test. Thus, the Kruskal-Wallis test revealed significant differences in the average duration of whole fixations among groups (H (2) = 7.489, p = .023). Similarly, the Mann-Whitney U test indicated significant differences in the number of saccades between positions (U = 40502.0, p = .035) and between expert and student groups (U = 14653.0, p = .029). Furthermore, the Kruskal-Wallis test showed notable differences in the number of saccades among groups (H (2) = 6.346, p= .042). All p-values were significant at the $\alpha$ = 0.05 level in this result.
@@ 4.2. Questionnaires & FGI
Expert→Expert 10 3.72 1.24
Expert→Student 9 3.52 1.15
Student→Expert 9 3.90 1.02
Student→Student 8 3.63 0.98